# Distributional Vision-Language Alignment by Cauchy-Schwarz Divergence

### v1

@@ 4.2 Image-Text Retrieval
Experimental Setup. Effective multimodal alignment also benefits cross-model retrieval. To demonstrate the alignment ability of our method on retrieval tasks, we conduct experiments aligning LLMs (Dubey et al., 2024) text representations with CLIP vision representations on both image-to-text and text-to-image retrieval. We use the Flickr 1K test set (Young et al., 2014) for short-text retrieval, while Urban1K (Zhang et al., 2025) and DOCCI (Onoe et al., 2025) are employed for long-text retrieval. We compare CS-Aligner against pure InfoNCE-based methods, such as Long-CLIP (Zhang et al., 2025) and LLM2CLIP (Huang et al., 2024), as the baselines. To ensure a fair comparison, we adopt the setup from LLM2CLIP, aligning CLIP ViT-L/14 image representations with Llama 3 (8B) text representations. Both the vision and text representations are aligned by adapters trained on the CC3M dataset.
@@
Comparisons. The results in Table 4 show that our method consistently and significantly outperforms the baselines across various datasets for both image-to-text (I2T) and text-to-image (T2I) retrieval. This demonstrates the effectiveness of our method for aligning the two modalities into a shared space. Moreover, the ability to align a different text encoder (LLM) with the successful alignment of an LLM-based text encoder with the CLIP image encoder highlights the flexibility and generalizability of our approach.
@@ 5 Conclusion
In this paper, we propose CS-Aligner, a novel distributional alignment framework that integrates Cauchy–Schwarz (CS) divergence with mutual information for multimodal alignment. By combining global distributional alignment with InfoNCE, CS-Aligner achieves tighter and more comprehensive alignment. By considering the modality distributional information, our method enables to leverage additional and detailed information from unpaired samples and tokens, leading to more flexible and fine-grained information for alignment. We demonstrate the effectiveness of our alignment on text-to-image generation and cross-modal retrieval.
Limitation & future work. Although our method has numerous applications, it is currently evaluated only on unCLIP-type models for generation. In the future, we will explore its integration with stable-diffusion-based models. Broader modalities (e.g., audio and video) and diverse tasks (e.g., image-to-text generation) could further strengthen alignment and broaden the applicability of CS-Aligner.
Impact Statement
This paper contributes to the advancement of Machine Learning. While our work may have various societal implications, none require specific emphasis at this stage.
Appendix A Details of the toy examples
@@
For Fig. 2 and Fig. 2, we set $\sigma_{\mathbf{x}}=\sigma_{\mathbf{y}}=1$. Hence, when $\mu_{\mathbf{x}}=\mu_{\mathbf{y}}=0$, $D_{\mathrm{KL}}\bigl{(}p(\mathbf{x})\,\|\,p(\mathbf{y})\bigr{)}=0$.
@@
For Fig. 2, we use $\sigma_{x}=10$ and $\sigma_{y}=0.1$. When $\mu_{\mathbf{x}}=0$ and $\mu_{\mathbf{y}}=2$, the $D_{\mathrm{KL}}\bigl{(}p(\mathbf{x})\,\|\,p(\mathbf{y})\bigr{)}=5194.89$, which is very large.
Appendix B Comparison bwtween CS divergence and other metrics
In the following, we analyze why CS divergence is a better choice than well-known KL divergence and Wasserstein distance.
@@ Appendix B Comparison bwtween CS divergence and other metrics
Comparison with KL divergence. KL divergence, a widely used metric in deep learning, is defined as:
$D_{\text{KL}}(p(\mathbf{x})\|p(\mathbf{y}))=\int p(\mathbf{x})\log\frac{p(\mathbf{x})}{p(\mathbf{y})}d\mathbf{x},$ (16)
where $p(\mathbf{x})$ and $p(\mathbf{y})$ are the probability distributions being compared. However, KL divergence is a non-symmetric metric, making it less suitable for learning a shared representation space (bi-directional), as commonly needed in multi-modal learning.
Minimal distribution overlap requirement. Furthermore, KL divergence can diverge to infinity when the distributions $p(\mathbf{x})$ and $p(\mathbf{y})$ have minimal overlap, making it invalid. In contrast, CS divergence remains well-defined under such conditions and provides a stable estimation (Yu et al., 2024; Yin et al., 2024). This issue is particularly critical in multi-modality settings, where the distributions from different modalities may have limited overlap.
@@
Nonparametric estimation. Additionally, when the distributions are not assumed to be Gaussian, a nonparametric estimator is required for KL divergence. A common choice, the k-NN estimator (Wang et al., 2009), is non-differentiable, which poses challenges for optimization in gradient-based learning frameworks. In contrast, the CS divergence demonstrates greater stability and differentiability when paired with KDE, making it a more robust choice.
@@
Comparison with wasserstein distance. Wasserstein Distance is also widely used for distribution discrepancy (e.g. GAN (Arjovsky et al., 2017)). However, Wasserstein distance is be computed either by using an additional learnable module (e.g., a neural network for estimating a transport map (Korotin et al., 2022)) or by solving an optimization problem, often approximated via multiple Sinkhorn (Cuturi, 2013) iterations for computational efficiency, leading to efficiency problem in large-scale training. In contrast, CS divergence can be efficiently estimated by a nonparametric estimator.
Appendix C Implementation details
@@
A proper kernel size is critical in KDE for accurate estimation of Eq. (8). In this paper, we follow (Yin et al., 2024) to normalize the features from two modalities and use a kernel size $1$. In general, this is sufficient to ensure stable learning.
C.1 T2I details

### v2

@@ 4.2 Image-Text Retrieval
Experimental Setup. Effective multimodal alignment also benefits cross-model retrieval. To demonstrate the alignment ability of our method on retrieval tasks, we align LLMs [25] text representations with CLIP vision representations on both image-to-text and text-to-image retrieval. We use the Flickr 1K test set [33] for short-text retrieval, while Urban1K [34] and DOCCI [35] are employed for long-text retrieval. We compare CS-Aligner against pure InfoNCE-based methods, such as Long-CLIP [34] and LLM2CLIP [1], as the baselines. To ensure a fair comparison, we adopt the setup from LLM2CLIP, aligning CLIP ViT-L/14 image representations with Llama 3 (8B) text representations. Both the vision and text representations are aligned by adapters trained on the CC3M dataset.
@@
Comparisons. Table 4 shows that our method consistently and significantly outperforms the baselines across various datasets for both image-to-text (I2T) and text-to-image (T2I) retrieval. This demonstrates the effectiveness of our method for aligning two modalities into a shared space. Moreover, the ability to align a different text encoder (LLM) with the CLIP image encoder highlights the flexibility and generalizability of our approach.
5 Related work
Vision-language alignment and applications. CLIP [6] serves as a foundational model for vision-language alignment in multimodal tasks. Several works have enhanced CLIP through techniques such as momentum distillation [36] and noisy text supervision [37]. Despite its success, CLIP suffers from a persistent modality gap between text and image representations. Prior studies [10, 11, 12] attribute this gap to factors such as cone effects [11] and suboptimal latent space structures [12]. To address this, various strategies have been proposed, including projection adapters [10, 13, 1], geodesic multimodal mixup [14], and parameter-efficient fine-tuning [38]. Recent works also improve CLIP by large language models (LLMs) [39, 2, 1] for downstream tasks such as image-text retrieval.
In addition to image-text retrieval, text-to-image (T2I) generation is another application that reflects the vision-language alignment capability. T2I has advanced significantly over the past decades, driven by both diffusion-based [40, 41, 42, 43] and GAN-based models [44, 45]. Among diffusion-based methods, the unCLIP framework [40, 3] employs a two-stage architecture with a CLIP-guided diffusion prior and a decoder (e.g., DALL-E-2 [3] or Karlo [31]). Its prior module $g_{\phi}$ maps text representations $\mathbf{y}$ to image ones $\mathbf{x}$ by a diffusion model. Recently, Eclipse [15] employs an $\ell_{2}$ loss to simplify the prior loss by eliminating diffusion time and intruding a noise $\epsilon$ term: $\mathcal{L}_{\text{prior}}=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}\left[\left\|\mathbf{x}-g_{\phi}(\epsilon,\mathbf{y})\right\|_{2}^{2}\right].$ However, these methods still rely on pairwise loss (e.g., $\ell_{2}$). In contrast, our approach introduces distributional alignment for a more holistic modality alignment.
Cauchy-Schwarz (CS) divergence. CS divergence [23, 18] is derived from the Cauchy-Schwarz inequality for square-integrable functions. It serves as a symmetric distribution distance metric with notable properties, such as the ability to measure conditional distributions [46] and the closed-form expression for mixtures of Gaussians [47]. CS divergence has been successfully applied across various domains, including deep clustering [48], disentangled representation learning [49], and deep regression [50]. Moreover, due to its advantage of estimating discrepancy between conditional distributions, it has demonstrated success in the domain adaption area [51] and time series clustering [46]. However, the utility of CS divergence in foundation models remains unclear and unexplored.
@@ 6 Conclusion
In this paper, we propose CS-Aligner, a novel distributional alignment framework that integrates Cauchy–Schwarz (CS) divergence with mutual information for multimodal alignment, which addresses the alignment and uniformity conflict of InfoNCE. By combining global distributional alignment with InfoNCE, CS-Aligner achieves tighter and more comprehensive alignment. By considering the modality distributional information, our method enables to leverage additional and detailed information from unpaired samples and tokens, leading to more flexible and fine-grained information for alignment. We demonstrate the effectiveness of our alignment on text-to-image generation and cross-modal retrieval.
Limitation & future work. Though our method has numerous applications, it is currently evaluated only on unCLIP-type models for generation. In future, we will explore its integration with stable-diffusion-based models. Broader modalities (e.g., audio and video) and diverse tasks (e.g., image-to-text generation) could further strengthen alignment and broaden the applicability of CS-Aligner.
Appendix A Broader Impacts
Our proposed method advances research in multimodal alignment by introducing a novel distributional alignment approach. As a result, it also facilitates progress in multimodal generation. In the meantime, this capability may raise ethical concerns, including the potential misuse for generating deceptive or inappropriate content.
Appendix B Details of the toy examples
Example B.1.
Consider two Gaussian distributions, $p(\mathbf{x})\sim\mathcal{N}(\mu_{\mathbf{x}},\sigma_{\mathbf{x}}^{2})$ and $p(\mathbf{y})\sim\mathcal{N}(\mu_{\mathbf{y}},\sigma_{\mathbf{y}}^{2})$, with a joint distribution $p(\mathbf{x},\mathbf{y})\sim\mathcal{N}\left(\begin{pmatrix}\mu_{\mathbf{x}}\\ \mu_{\mathbf{y}}\end{pmatrix},\begin{pmatrix}\sigma_{\mathbf{x}}^{2}&\rho\sigma_{\mathbf{x}}\sigma_{\mathbf{y}}\\ \rho\sigma_{\mathbf{x}}\sigma_{\mathbf{y}}&\sigma_{\mathbf{y}}^{2}\end{pmatrix}\right)$. Here, $\mu_{\mathbf{x}}$ and $\mu_{\mathbf{y}}$ are the means of $\mathbf{x}$ and $\mathbf{y}$, $\sigma_{\mathbf{x}}^{2}$ and $\sigma_{\mathbf{y}}^{2}$ are their variances, and $\rho$ is the correlation coefficient and controls their linear dependency. When $\rho=0.99$, the two modalities are highly dependent, with high mutual information ($I=1.959$; see Fig. 2a and 2b). When $\rho=0$, the modalities are independent, resulting in zero mutual information (Fig. 2c). Interestingly, two distributions with the same mutual information value can either exhibit minimal statistical distance and nearly identical shapes, including similar locations, widths, and higher-order moments, as shown in Fig. 2b, or have completely different shapes with distinct means (00 for $p(\mathbf{x})$ and $2$ for $p(\mathbf{y})$) and variances ($4$ for $p(\mathbf{x})$ and $1$ for $p(\mathbf{y})$), as illustrated in Fig. 2a. Quantitatively, the former case shows a minimal KL divergence of 00, while the latter exhibits a KL divergence of nearly $6.81$.
@@ Appendix B Details of the toy examples
For Fig. 2b and Fig. 2c, we set $\sigma_{\mathbf{x}}=\sigma_{\mathbf{y}}=1$. Hence, when $\mu_{\mathbf{x}}=\mu_{\mathbf{y}}=0$, $D_{\mathrm{KL}}\bigl{(}p(\mathbf{x})\,\|\,p(\mathbf{y})\bigr{)}=0$.
@@
For Fig. 2a, we use $\sigma_{x}=2$ and $\sigma_{y}=1$. When $\mu_{\mathbf{x}}=0$ and $\mu_{\mathbf{y}}=2$, the $D_{\mathrm{KL}}\bigl{(}p(\mathbf{x})\,\|\,p(\mathbf{y})\bigr{)}\approx 6.81$, which is very large.
Appendix C Derivations
In this section, we provide a derivation of alignment and uniformity terms of InfoNCE. More concrete analysis can be found in [22].
Let $(\mathbf{x},\mathbf{y})$ be positive (image–text) pairs drawn from $p_{\text{pair}}$, and let $\{(\mathbf{x}^{\prime}_{i},\mathbf{y}^{\prime}_{i})\}_{i=1}^{M}$ be $M$ negative samples (unpaired samples) drawn i.i.d. from the marginal $p_{\text{data}}$. The one-sided InfoNCE (CLIP) loss with temperature $\tau>0$ is
$\mathcal{L}_{\text{InfoNCE}}=-\frac{1}{2}\,\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p_{\text{pair}}}\mathbb{E}_{\{\mathbf{x}^{\prime}_{i},\mathbf{y}^{\prime}_{i}\}\sim p_{\text{data}}}\Biggl{[}\log\frac{e^{\,\mathbf{x}^{\top}\mathbf{y}/\tau}}{\sum_{i=1}^{M}e^{\,\mathbf{x}_{i}^{\prime}{}^{\!T}\mathbf{y}/\tau}}+\,\log\frac{e^{\,\mathbf{x}^{\top}\mathbf{y}/\tau}}{\sum_{i=1}^{M}e^{\,\mathbf{x}^{\top}\mathbf{y}^{\prime}_{i}/\tau}}\Biggr{]}.$
In CLIP, the features are normalized to compute the loss. Under this unit-norm constraint $\|\mathbf{x}\|_{2}=\|\mathbf{y}\|_{2}=1$, $\mathcal{L}_{\text{InfoNCE}}$ decomposes into
$\mathcal{L}_{\text{InfoNCE}}=\underbrace{-\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p_{\mathrm{pair}}}\!\Bigl{[}\frac{\mathbf{x}^{\top}\mathbf{y}}{\tau}\Bigr{]}}_{\mathcal{L}_{\text{align}}}\;+\;\underbrace{\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p_{\text{data}}}\Bigl{[}\tfrac{1}{2}\log\!\sum_{i=1}^{M}e^{\,\mathbf{x}^{\top}\mathbf{y}^{\prime}_{i}/\tau}+\tfrac{1}{2}\log\!\sum_{i=1}^{M}e^{\,\mathbf{x}^{\prime}_{i}{}^{\!T}\mathbf{y}/\tau}\Bigr{]}}_{\mathcal{L}_{\mathrm{uniform}}},$
up to an additive constant. Moreover, by writing
$\|\mathbf{x}-\mathbf{y}\|_{2}^{2}=\|\mathbf{x}\|^{2}+\|y\|^{2}-2\,\mathbf{x}^{\top}\mathbf{y}=2-2\,\mathbf{x}^{\top}\mathbf{y}\quad\Longrightarrow\quad\mathbf{x}^{\top}\mathbf{y}=1-\tfrac{1}{2}\|\mathbf{x}-\mathbf{y}\|_{2}^{2},$ (16)
we can show that:
  1. (i)
    Alignment.
    $-\,\mathbb{E}_{(\mathbf{x},\mathbf{y})}\Bigl{[}\frac{\mathbf{x}^{\top}\mathbf{y}}{\tau}\Bigr{]}$ $=-\,\mathbb{E}\!\Bigl{[}\frac{1-\tfrac{1}{2}\|\mathbf{x}-\mathbf{y}\|^{2}}{\tau}\Bigr{]}=-\frac{1}{\tau}+\frac{1}{2\tau}\,\mathbb{E}\bigl{[}\|\mathbf{x}-\mathbf{y}\|^{2}\bigr{]}.$
    Dropping the constant $-1/\tau$, define
    $\mathcal{L}_{\text{align}}:=\frac{1}{2\tau}\,\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p_{\text{pair}}}\bigl{[}\|\mathbf{x}-\mathbf{y}\|_{2}^{2}\bigr{]}.$ (17)
  2. (ii)
    Uniformity. For each negative (unpaired) sample $y^{\prime}_{i}$, using Eq. (16),
    $e^{\,\mathbf{x}^{\top}\mathbf{y}^{\prime}_{i}/\tau}=e^{(1-\tfrac{1}{2}\|\mathbf{x}-\mathbf{y}^{\prime}_{i}\|^{2})/\tau}=e^{1/\tau}\,e^{-\tfrac{1}{2\tau}\|\mathbf{x}-\mathbf{y}^{\prime}_{i}\|^{2}}.$
    Hence
    $\sum_{i=1}^{M}e^{\,\mathbf{x}^{\top}\mathbf{y}^{\prime}_{i}/\tau}=e^{1/\tau}\sum_{i=1}^{M}e^{-\tfrac{1}{2\tau}\|\mathbf{x}-\mathbf{y}^{\prime}_{i}\|^{2}},\quad\log\!\sum_{i}e^{\,\mathbf{x}^{\top}\mathbf{y}^{\prime}_{i}/\tau}=\tfrac{1}{\tau}+\log\!\sum_{i}e^{-\tfrac{1}{2\tau}\|\mathbf{x}-\mathbf{y}^{\prime}_{i}\|^{2}}.$
    An identical argument holds for the $\{\mathbf{x}^{\prime}_{i},\mathbf{y}\}$ terms. Up to constants,
    $\mathcal{L}_{\mathrm{uniform}}:=\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p_{\mathrm{data}}}\Bigl{[}\tfrac{1}{2}\log\!\sum_{i=1}^{M}e^{-\tfrac{1}{2\tau}\|\mathbf{x}-\mathbf{y}^{\prime}_{i}\|^{2}}+\tfrac{1}{2}\log\!\sum_{i=1}^{M}e^{-\tfrac{1}{2\tau}\|\mathbf{x}^{\prime}_{i}-\mathbf{y}\|^{2}}\Bigr{]}.$ (18)
In the limit of large batch size one may further rewrite
$\mathcal{L}_{\mathrm{uniform}}\;\approx\;\log\,\mathbb{E}_{\mathbf{x},\mathbf{y}\sim p_{\mathrm{data}}}\bigl{[}\exp(-t\|\mathbf{x}-\mathbf{y}\|_{2}^{2})\bigr{]},$
with $t=\tfrac{1}{2\tau}$.
Combining (i) and (ii) and absorbing all additive constants gives the desired decomposition
$\boxed{\mathcal{L}_{\mathrm{clip}}=\mathcal{L}_{\mathrm{align}}+\mathcal{L}_{\mathrm{uniform}}+\text{const.}}$.
Appendix D Comparison between CS divergence and other metrics
Unlike parametric distributions, distributions of different real-world modalities exhibit unpredictable variability and inconsistent overlaps, meaning that $p(\mathbf{x})$ and $p(\mathbf{y})$ may follow arbitrary distributions with a small intersection. Therefore, it is crucial to overcome these challenges to measure and optimize multimodal distribution divergence robustly. Below, we outline several key properties that an effective metric should satisfy for multimodal alignment.
Remark D.1.
Key properties for distribution align metrics:
These properties enable the divergence term to align arbitrary distributions with small support overlap, which is well-suited for large-scale multimodal applications involving deep learning.
D.1 Connection to the prior loss
Remark D.2.
Connection to the prior loss ($\ell_{2}$ loss) used by Eclipse [15]:
Consider the third term in Eq. (8), which involves $\kappa(\mathbf{x}_{i},\mathbf{y}_{j})$ defined by the Gaussian kernel $\kappa_{\sigma}(\mathbf{x},\mathbf{y})=\exp\!\bigl{(}-\|\mathbf{x}-\mathbf{y}\|_{2}^{2}/2\sigma^{2}\bigr{)}$. A second-order Taylor expansion yields
D.2 Comparison with KL divergence.
@@ D.1 Connection to the prior loss
KL divergence is a widely used metric in deep learning. Given two distributions, $p(\omega)$ and $q(\omega)$, the KL divergence is defined as:
$D_{\mathrm{KL}}(p;q)=\int p(\omega)\log\frac{p(\omega)}{q(\omega)}\,d\omega$ (21)
Validity for multimodal alignment. Define the support sets of distributions $p$ and $q$ as:
$\text{supp}(p)=\{\omega\in\Omega:p(\omega)>0\},\quad\text{supp}(q)=\{\omega\in{\Omega}:q(\omega)>0\}.$ (22)
For KL divergence, if there exists any point $x\in\text{supp}(p)$ such that $q(x)=0$, the term $p(\omega)\log\frac{p(\omega)}{0}\rightarrow\infty$, leading to: $D_{\mathrm{KL}}(p;q)=\infty.$ Thus, a necessary condition for KL divergence to be finite is $\text{supp}(p)\subseteq\text{supp}(q).$ Otherwise, KL divergence becomes invalid.
[…]
In multimodal alignment, it’s reasonable to assume that the two modality distributions partially overlap but are not disjoint, as supported by our empirical observations in Fig. 4a (no alignment results). Under these conditions, KL divergence can be invalid and therefore suboptimal. Conversely, the CS divergence condition is less restrictive, making it more suitable and stable for multimodal alignment.
Compatibility with InfoNCE Integrating InfoNCE with CS divergence explicitly encourages intra-modality uniformity and cross-modality alignment, thereby effectively improving multimodal alignment. For KL divergence, assuming the distributions of the two modalities are Gaussian, $\mathcal{N}(\mu_{0},\Sigma_{0})$ and $\mathcal{N}(\mu_{1},\Sigma_{1})$, the divergence can be computed as:
$\mathcal{D}_{\mathrm{KL}}\left[\mathcal{N}(\mu_{0},\Sigma_{0})\|\mathcal{N}(\mu_{1},\Sigma_{1})\right]=\frac{1}{2}\left(\mathrm{tr}\left(\Sigma_{1}^{-1}\Sigma_{0}\right)+(\mu_{1}-\mu_{0})^{\top}\Sigma_{1}^{-1}(\mu_{1}-\mu_{0})-k+\log\left(\frac{\det\Sigma_{1}}{\det\Sigma_{0}}\right)\right).$ (23)
This formulation lacks explicit connections to the InfoNCE in terms of alignment and uniformity, making it less compatible with InfoNCE compared to the CS divergence.
@@
Nonparametric estimation. Additionally, when the distributions are not assumed to be Gaussian, a nonparametric estimator is required for KL divergence. A common choice, the k-NN estimator [52], is non-differentiable, which poses challenges for optimization in gradient-based learning frameworks. In contrast, the CS divergence demonstrates greater stability and differentiability when paired with KDE, making it a more robust choice.
Experimental Comparison. To verify the above analysis, we compare CS divergence and KL divergence on the unpaired data scenario, where KL can easily become invalid. We trained a KL + InfoNCE model in our unpaired data setting—using paired data for InfoNCE and unpaired data for divergence. The initial KL value exceeded $5000$ (extremely large), and consequently, the model could not converge, leading to catastrophic failure. In contrast, CS divergence remained stable (initial value around 3), and achieved comparable final performance with an FID of 12.18 (Fig. 5(b) in the main paper).
For the scenario
D.3 Comparison with Wasserstein distance.
@@
Wasserstein Distance is also widely used for distribution discrepancy (e.g. GAN [53]). However, Wasserstein distance is be computed either by using an additional learnable module (e.g., a neural network for estimating a transport map [54]) or by solving an optimization problem, often approximated via multiple Sinkhorn [55] iterations for computational efficiency, leading to efficiency problem in large-scale training. In contrast, CS divergence can be efficiently estimated by a nonparametric estimator.
D.4 Comparison with mutual information divergence [56].
Mutual information estimation depends on parametric assumptions about the underlying distributions—e.g., multivariate Gaussian—whereas CS divergence imposes no such constraints. Moreover, estimating mutual information decomposes into a mutual information term plus two KL divergences, and thus lacks explicit connections to the InfoNCE in terms of alignment and uniformity.
Appendix E More Results
E.1 More Visualization
E.2 More visualization for token alignment
Appendix F Implementation details
@@
A proper kernel size is critical in KDE for accurate estimation of Eq. (8). In this paper, we follow [51] to normalize the features from two modalities and use a kernel size $1$. In general, this is sufficient to ensure stable learning.
F.1 T2I details